# What Factors Affect LLMs and RLLMs in Financial Question Answering?

### v1

@@ Abstract
Recently, the development of large language models (LLMs) and reasoning large language models (RLLMs) have gained considerable attention from many researchers. RLLMs enhance the reasoning capabilities of LLMs through Long Chain-of-Thought (Long CoT) processes, significantly improving the performance of LLMs in addressing complex problems. However, there are few works that systematically explore what methods can fully unlock the performance of LLMs and RLLMs within the financial domain. To investigate the impact of various methods on LLMs and RLLMs, we utilize five LLMs and three RLLMs to assess the effects of prompting methods, agentic frameworks, and multilingual alignment methods on financial question-answering tasks. Our research findings indicate: (1) Current prompting methods and agent frameworks enhance the performance of LLMs in financial question answering by simulating Long CoT; (2) RLLMs possess inherent Long CoT capabilities, which limits the effectiveness of conventional methods in further enhancing their performance; (3) Current advanced multilingual alignment methods primarily improve the multilingual performance of LLMs by extending the reasoning length, which yields minimal benefits for RLLMs. We hope that this study can serve as an important reference for LLMs and RLLMs in the field of financial question answering.
What Factors Affect LLMs and RLLMs in Financial Question Answering?
[…]
@@ 1 Introduction
Recently, large language models (LLMs) have significantly advanced the field of natural language processing (NLP), and more and more researchers utilize LLMs to solve complex tasks in various domains Zhou et al. (2023); Nie et al. (2024); Qin et al. (2024); Chang et al. (2024); Wang et al. (2025a). Furthermore, some researchers propose prompting-based and agentic frameworks to enhance the capabilities of LLMs Wei et al. (2022); Yao et al. (2023a); Zhang et al. (2024b); Yao et al. (2023b); Li et al. (2023); Hong et al. (2024), which further expands the capabilities of LLMs in tasks in different fields. Particularly, to enhance the reasoning capabilities of LLMs in addressing complex problems, researchers introduce reasoning large language models (RLLMs). The application of the Long Chain-of-Thought (Long CoT) fully unlocks the understanding ability of RLLMs Chen et al. (2025). Long CoT entails deeper reasoning, reflective analysis, and a more extensive exploration of logical structures Gandhi et al. (2025), which enables RLLMs to perform on par with experts in certain domains.
Researchers have explored the application of LLMs in financial question answering. Wu et al. (2023) present BloombergGPT, a 50 billion parameter LLM for financial domain. Fatemi and Hu (2024) propose a multi-agent framework to enhance the performance of LLMs in financial question answering. Xie et al. (2024) introduce a novel financial benchmark, FinBen, including 42 datasets spanning 24 financial tasks. Srivastava et al. (2024) evaluate LLMs in four financial tabular question-answering datasets. Xue et al. (2024) build a benchmark for financial multilingual multi-modal question answering named FAMMA, which is a challenging benchmark for multi-modal LLMs. Wang et al. (2025b) propose FinSage, which solves the financial filings question answering task through retrieval augmented generation (RAG).
Although LLMs have made significant progress in the field of financial question answering, existing work still mainly focuses on enhancing their performance within this domain. With the advancement of LLMs and RLLMs, the factors influencing their financial question-answering capabilities remain under-explored. Inspired by this gap, we systematically investigate the following issues: ($i$) the impact of prompting methods on LLMs/RLLMs; ($ii$) the influence of agentic frameworks on LLMs/RLLMs; and ($iii$) whether multilingual alignment methods can enhance the multilingual financial question-answering abilities of LLMs/RLLMs. We utilize five LLMs and three RLLMs, conducting detailed experiments across seven representative methods. Our code will be available at https://github.com/WPENGxs/LLM_RLLM_financial_analysis.
@@ 3.1 Exploration of prompting methods
Current research extensively demonstrates that different prompting methods can significantly alter the performance of LLMs Wei et al. (2022); Wang et al. (2023); Zhang et al. (2024b). The prompting method modifies certain aspects of the thinking prompt while ensuring the question remains intact, enabling LLMs to solve the question by following the structured steps of the thinking prompt.
We select three representative prompting methods to investigate the impact of varying prompt word changes on the performance of LLMs and RLLMs: Direct, Zero-shot CoT Kojima et al. (2022), and Plan-and-Solve Wang et al. (2023). These prompting methods can be defined as:
@@ 3.2 Exploration of agentic frameworks
To further enhance the performance of LLMs, researchers have started to incorporate agentic systems. Specifically, by enabling LLMs to assume different roles and engage in continuous communication and cooperation, their performance can be significantly improved Luo et al. (2025).
We select two agentic frameworks, and aim to explore the impact of LLMs and RLLMs in different agent collaborations: Self-Refine Madaan et al. (2023) and S3 Agent Wang et al. (2024). These agentic frameworks can be defined as:
@@
  • S3 Agent: S3 Agent proposes three agents of different perspectives: superficial expression, semantic information, and sentiment expression, to unlock the power of LLMs.
@@ 3.3 Exploration of multilingual alignment methods
We evaluate the performance of four standard multilingual approaches, including state-of-the-art cross-lingual prompting methods, on multilingual alignment: Direct, Translate-en Shi et al. (2023), and Cross-lingual Prompting Qin et al. (2023).
@@
Method Arithmetic Non-Arithmetic Overall Overall Easy Medium Hard Overall Easy Medium Hard Llama-3.1-8B-Instruct Grattafiori et al. (2024) Direct Grattafiori et al. (2024) 7.53 9.17 5.65 6.88 14.36 14.83 14.73 13.96 10.59 Zero-shot CoT Kojima et al. (2022) 8.18 11.35 4.24 7.18 22.18 23.30 20.00 22.52 14.44 Plan-and-Solve Wang et al. (2023) 11.16 13.53 9.18 9.58 22.06 25.84 17.89 21.84 16.04 Self-Refine Madaan et al. (2023) 13.67 18.55 8.12 11.67 25.28 26.27 23.68 25.45 18.86 S3 Agent Wang et al. (2024) 14.97 17.90 9.54 15.56 21.95 28.38 18.94 19.81 18.09 GPT-4o-mini Hurst et al. (2024) Direct Hurst et al. (2024) 22.32 30.56 18.02 14.67 30.80 37.71 29.47 27.70 26.11 Zero-shot CoT Kojima et al. (2022) 24.09 29.91 21.20 18.56 33.67 44.91 30.00 29.27 28.38 Plan-and-Solve Wang et al. (2023) 24.93 31.87 21.55 18.26 33.67 42.79 29.47 30.63 28.84 Self-Refine Madaan et al. (2023) 22.23 28.38 18.72 16.76 32.18 41.10 27.89 29.27 26.68 S3 Agent Wang et al. (2024) 21.86 28.82 18.37 15.26 31.49 41.10 28.42 27.70 26.16 Gemini-1.5-flash Team et al. (2024) Direct Team et al. (2024) 25.02 33.18 19.08 18.86 32.29 42.79 26.84 29.05 28.27 Zero-shot CoT Kojima et al. (2022) 24.74 32.53 19.08 18.86 32.52 42.79 27.89 29.05 28.22 Plan-and-Solve Wang et al. (2023) 25.76 33.18 22.61 18.26 33.33 42.37 28.94 30.40 29.15 Self-Refine Madaan et al. (2023) 21.86 28.16 16.60 17.66 29.65 36.44 25.78 27.70 25.34 S3 Agent Wang et al. (2024) 25.95 33.40 20.14 20.65 32.52 41.52 30.00 28.82 28.89 Qwen-2.5-32B Yang et al. (2024) Direct Yang et al. (2024) 27.90 35.80 22.61 21.55 35.17 43.64 31.57 32.20 31.15 Zero-shot CoT Kojima et al. (2022) 27.62 34.49 23.67 21.55 35.28 41.52 31.05 33.78 31.05 Plan-and-Solve Wang et al. (2023) 24.83 30.34 23.32 18.56 34.94 44.49 31.05 31.53 29.35 Self-Refine Madaan et al. (2023) 27.44 33.84 24.38 21.25 36.43 41.94 30.52 36.03 31.46 S3 Agent Wang et al. (2024) 29.20 36.68 24.38 23.05 35.28 43.64 34.73 31.08 31.92 DeepSeek-V3 Liu et al. (2024) Direct Liu et al. (2024) 35.06 42.13 29.32 30.23 43.21 51.27 35.78 42.11 38.71 Zero-shot CoT Kojima et al. (2022) 36.09 43.23 31.44 30.23 43.67 51.69 35.78 42.79 39.48 Plan-and-Solve Wang et al. (2023) 35.53 43.01 29.32 30.53 44.94 51.27 37.36 44.81 39.74 Self-Refine Madaan et al. (2023) 33.58 41.04 27.91 28.14 41.49 47.88 34.21 41.21 37.12 S3 Agent Wang et al. (2024) 35.34 41.92 30.38 30.53 42.52 51.27 35.78 40.76 38.56 DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025) Direct Guo et al. (2025) 33.67 42.79 27.20 26.64 38.73 47.03 31.57 37.38 35.93 Zero-shot CoT Kojima et al. (2022) 34.69 42.79 30.03 27.54 38.27 49.15 30.52 35.81 36.29 Plan-and-Solve Wang et al. (2023) 33.86 42.79 26.85 27.54 36.20 45.76 30.52 33.55 34.91 Self-Refine Madaan et al. (2023) 28.37 35.15 24.02 22.75 33.79 38.13 27.36 34.23 30.79 S3 Agent Wang et al. (2024) 34.41 42.13 29.68 27.84 39.65 50.42 32.10 37.16 36.76 Qwen-3-14B Yang et al. (2025) Direct Yang et al. (2025) 33.30 43.23 28.97 23.35 38.62 49.57 32.63 35.36 35.68 Zero-shot CoT Kojima et al. (2022) 34.23 43.66 28.62 26.04 38.50 47.88 32.63 36.03 36.14 Plan-and-Solve Wang et al. (2023) 33.58 44.10 28.97 23.05 37.12 49.15 30.52 33.55 35.16 Self-Refine Madaan et al. (2023) 33.48 43.23 28.97 23.95 40.57 50.42 35.26 37.61 36.65 S3 Agent Wang et al. (2024) 34.04 43.44 30.03 24.55 40.68 51.69 33.68 37.83 37.01 O4-mini OpenAI (2025) Direct OpenAI (2025) 39.25 47.16 34.62 32.33 49.65 54.66 39.47 51.35 43.90 Zero-shot CoT Kojima et al. (2022) 39.62 46.50 33.21 35.62 50.57 54.23 42.63 52.01 44.52 Plan-and-Solve Wang et al. (2023) 39.72 46.50 34.27 35.02 49.54 53.81 40.52 51.12 44.11 Self-Refine Madaan et al. (2023) 38.97 46.06 31.44 35.62 49.08 52.11 39.47 51.57 43.49 S3 Agent Wang et al. (2024) 39.81 46.72 33.21 35.92 50.11 54.66 38.42 52.70 44.42
@@ 4.1 Experimental Settings
We use the standard financial question answering benchmark FAMMA Xue et al. (2024) for experiments. FAMMA is a benchmark for financial multilingual multi-modal question answering, which includes English, Chinese, and French. Since current RLLMs are unimodal and cannot process multimodal information, we utilize the Basic Txt dataset in FAMMA, which converts multimodal data into textual information using OCR. The final dataset used for evaluation comprises 1,945 entries.
We conduct experiments on seven backbones, including five LLMs and three RLLMs: Meta-Llama-3.1-8B-Instruct Grattafiori et al. (2024), GPT-4o-mini Hurst et al. (2024), Gemini-1.5-flash Team et al. (2024), Qwen-2.5-32B Yang et al. (2024), DeepSeek-V3 Liu et al. (2024), DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025), thing mode of Qwen-3-14B Yang et al. (2025), O4-mini OpenAI (2025). All LLMs and RLLMs use default settings.
@@ 4.2 Analysis of Prompting Methods
Larger LLMs exhibit increased robustness to various prompting methods. It is evident that for larger LLMs, such as Gemini-1.5-flash and DeepSeek-V3, the performance gap between different prompting methods is not significant, with performance fluctuations remaining under 1%. This indicates that LLMs with superior performance are less affected by input prompts, demonstrating greater robustness to variations in input.
Method Chinese French Overall (Chinese+French) Meta-Llama-3.1-8B-Instruct Grattafiori et al. (2024) Direct Grattafiori et al. (2024) 11.43 5.12 9.70 Translate-en Shi et al. (2023) 12.40 (+0.97) 6.41 (+1.29) 10.75 (+1.05) CLP Qin et al. (2023) 15.57 (+4.14) 5.12 (+0.00) 12.69 (+2.99) GPT-4o-mini Hurst et al. (2024) Direct Hurst et al. (2024) 24.81 13.46 21.69 Translate-en Shi et al. (2023) 25.79 (+0.98) 14.10 (+0.64) 22.57 (+0.88) CLP Qin et al. (2023) 27.98 (+3.17) 12.17 (-1.29) 23.63 (+1.94) Gemini-1.5-flash Team et al. (2024) Direct Team et al. (2024) 27.25 10.25 22.57 Translate-en Shi et al. (2023) 27.00 (-0.25) 9.61 (-0.64) 22.22 (-0.35) CLP Qin et al. (2023) 27.73 (+0.48) 9.61 (-0.64) 22.75 (+0.18) Qwen-2.5-32B Yang et al. (2024) Direct Yang et al. (2024) 30.17 13.46 25.57 Translate-en Shi et al. (2023) 32.36 (+2.19) 13.46 (+0.00) 27.16 (+1.59) CLP Qin et al. (2023) 31.14 (+0.97) 14.10 (+0.64) 26.45 (+0.88) DeepSeek-V3 Liu et al. (2024) Direct Liu et al. (2024) 38.44 22.43 34.03 Translate-en Shi et al. (2023) 38.44 (+0.00) 23.07 (+0.64) 34.21 (+0.18) CLP Qin et al. (2023) 40.87 (+2.43) 25.64 (+3.21) 36.68 (+2.65) DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025) Direct Guo et al. (2025) 34.06 23.71 31.21 Translate-en Shi et al. (2023) 35.52 (+1.46) 21.15 (-2.56) 31.56 (+0.35) CLP Qin et al. (2023) 35.52 (+1.46) 19.23 (-4.48) 31.04 (-0.17) Qwen-3-14B Yang et al. (2025) Direct Yang et al. (2025) 37.22 19.87 32.45 Translate-en Shi et al. (2023) 38.19 (+0.97) 21.15 (+1.28) 33.50 (+1.05) CLP Qin et al. (2023) 35.76 (-1.46) 23.07 (+3.20) 32.27 (-0.18) O4-mini OpenAI (2025) Direct OpenAI (2025) 42.82 26.92 38.44 Translate-en Shi et al. (2023) 44.28 (+1.46) 26.28 (-0.64) 39.32 (+0.88) CLP Qin et al. (2023) 43.76 (+0.94) 26.92 (+0.00) 39.11 (+0.67)

### v3

@@ Abstract
Recently, the development of large language models (LLMs) and reasoning large language models (RLLMs) have gained considerable attention from many researchers. RLLMs enhance the reasoning capabilities of LLMs through Long Chain-of-Thought (Long CoT) processes, significantly improving the performance of LLMs in addressing complex problems. However, there are few works that systematically explore what methods can fully unlock the performance of LLMs and RLLMs within the financial domain. To investigate the impact of various methods on LLMs and RLLMs, we utilize five LLMs and three RLLMs to assess the effects of prompting methods, agentic frameworks, and multilingual alignment methods on financial question-answering tasks. Our research findings indicate: (1) Current prompting methods and agent frameworks enhance the performance of LLMs in financial question answering by simulating Long CoT; (2) RLLMs possess inherent Long CoT capabilities, which limits the effectiveness of conventional methods in further enhancing their performance; (3) Current advanced multilingual alignment methods primarily improve the multilingual performance of LLMs by extending the reasoning length, which yields minimal benefits for RLLMs. Additionally, we discuss strategies for enhancing the performance of LLMs and RLLMs in financial question answering, which may serve as a inspiration for future improvements. We hope that this study can serve as an important reference for LLMs and RLLMs in the field of financial question answering.
@@ 1 Introduction
Recently, large language models (LLMs) have significantly advanced the field of natural language processing (NLP), and more and more researchers utilize LLMs to solve complex tasks in various domains (Zhou et al., 2023; Nie et al., 2024; Qin et al., 2024; Chang et al., 2024; Wang et al., 2025a). Furthermore, some researchers propose prompting-based and agentic frameworks to enhance the capabilities of LLMs (Wei et al., 2022; Yao et al., 2023a; Zhang et al., 2024b; Yao et al., 2023b; Li et al., 2023; Hong et al., 2024), which further expands the capabilities of LLMs in tasks in different fields. Particularly, to enhance the reasoning capabilities of LLMs in addressing complex problems, researchers introduce reasoning large language models (RLLMs). The application of the Long Chain-of-Thought (Long CoT) fully unlocks the understanding ability of RLLMs (Chen et al., 2025). Long CoT entails deeper reasoning, reflective analysis, and a more extensive exploration of logical structures (Gandhi et al., 2025), which enables RLLMs to perform on par with experts in certain domains.
Researchers have explored the application of LLMs in financial question answering. Wu et al. (2023) present BloombergGPT, a 50 billion parameter LLM for financial domain. Fatemi & Hu (2024) propose a multi-agent framework to enhance the performance of LLMs in financial question answering. Xie et al. (2024) introduce a novel financial benchmark, FinBen, including 42 datasets spanning 24 financial tasks. Srivastava et al. (2024) evaluate LLMs in four financial tabular question-answering datasets. Xue et al. (2024) build a benchmark for financial multilingual multi-modal question answering named FAMMA, which is a challenging benchmark for multi-modal LLMs. Wang et al. (2025c) propose FinSage, which solves the financial filings question answering task through retrieval augmented generation (RAG).
Although LLMs have made significant progress in the field of financial question answering, existing work still mainly focuses on enhancing their performance within this domain. With the advancement of LLMs and RLLMs, the factors influencing their financial question-answering capabilities remain under-explored. Inspired by this gap, we systematically investigate the following issues: ($i$) the impact of prompting methods on LLMs/RLLMs; ($ii$) the influence of agentic frameworks on LLMs/RLLMs; and (iiiiii) whether multilingual alignment methods can enhance the multilingual financial question-answering abilities of LLMs/RLLMs. We utilize five LLMs and three RLLMs, conducting detailed experiments across seven representative methods.
@@ 3.1 Exploration of prompting methods
Current research extensively demonstrates that different prompting methods can significantly alter the performance of LLMs (Wei et al., 2022; Wang et al., 2023; Zhang et al., 2024b). The prompting method modifies certain aspects of the thinking prompt while ensuring the question remains intact, enabling LLMs to solve the question by following the structured steps of the thinking prompt.
We select three representative prompting methods to investigate the impact of varying prompt word changes on the performance of LLMs and RLLMs: Direct, Zero-shot CoT (Kojima et al., 2022), and Plan-and-Solve (Wang et al., 2023). These prompting methods can be defined as:
@@ 3.2 Exploration of agentic frameworks
To further enhance the performance of LLMs, researchers have started to incorporate agentic systems. Specifically, by enabling LLMs to assume different roles and engage in continuous communication and cooperation, their performance can be significantly improved (Luo et al., 2025a).
We select two agentic frameworks, and aim to explore the impact of LLMs and RLLMs in different agent collaborations: Self-Refine (Madaan et al., 2023) and S3 Agent (Wang et al., 2024). These agentic frameworks can be defined as:
@@
  • S3 Agent: S3 Agent proposes three agents of different perspectives: superficial expression, semantic information, and sentiment expression, to unlock the power of LLMs. We made minor modifications to the S3 Agent for financial question-answering tasks while preserving the three core components intact.
@@ 3.3 Exploration of multilingual alignment methods
We evaluate the performance of four standard multilingual approaches, including state-of-the-art cross-lingual prompting methods, on multilingual alignment: Direct, Translate-en (Shi et al., 2023), and Cross-lingual Prompting (Qin et al., 2023).
@@
Method Arithmetic Non-Arithmetic Overall Overall Easy Medium Hard Overall Easy Medium Hard Llama-3.1-8B-Instruct (Grattafiori et al., 2024) Direct (Grattafiori et al., 2024) 10.88 12.66 9.89 9.28 23.44 22.45 32.10 20.27 16.50 Zero-shot CoT (Kojima et al., 2022) 13.20 16.15 10.95 11.07 31.72 32.62 34.73 29.95 21.49 Plan-and-Solve (Wang et al., 2023) 16.09 18.99 14.84 13.17 29.77 32.62 30.52 27.92 22.21 Self-Refine (Madaan et al., 2023) 17.95 22.70 13.07 15.56 32.06 34.74 35.26 29.27 24.26 S3 Agent (Wang et al., 2024) 19.62 22.27 15.54 19.46 30.80 37.28 32.63 26.57 24.62 GPT-4o-mini (Hurst et al., 2024) Direct (Hurst et al., 2024) 32.46 41.70 30.38 21.55 42.52 50.42 50.52 34.90 36.96 Zero-shot CoT (Kojima et al., 2022) 34.13 41.04 34.27 24.55 45.17 56.35 52.10 36.26 39.07 Plan-and-Solve (Wang et al., 2023) 36.18 44.97 33.92 26.04 45.74 56.77 51.05 37.61 40.46 Self-Refine (Madaan et al., 2023) 33.30 41.26 33.21 22.45 42.06 52.11 45.78 35.13 37.22 S3 Agent (Wang et al., 2024) 30.88 38.42 28.26 22.75 41.72 51.69 46.84 34.23 35.73 Gemini-1.5-flash (Team et al., 2024) Direct (Team et al., 2024) 37.20 47.16 35.33 25.14 45.28 58.05 49.47 36.71 40.82 Zero-shot CoT (Kojima et al., 2022) 37.39 46.72 36.04 25.74 46.09 58.89 51.57 36.93 41.28 Plan-and-Solve (Wang et al., 2023) 37.76 46.50 39.57 24.25 47.12 55.50 55.26 39.18 41.95 Self-Refine (Madaan et al., 2023) 34.41 42.35 33.92 23.95 42.52 52.96 50.00 33.78 38.04 S3 Agent (Wang et al., 2024) 37.20 45.85 34.62 27.54 44.59 57.20 50.52 35.36 40.51 Qwen-2.5-32B (Yang et al., 2024) Direct (Yang et al., 2024) 41.86 50.65 40.98 30.53 48.62 59.32 56.31 39.63 44.88 Zero-shot CoT (Kojima et al., 2022) 42.32 50.21 40.98 32.63 50.80 59.32 60.00 42.34 46.11 Plan-and-Solve (Wang et al., 2023) 39.44 44.97 43.46 28.44 49.77 60.59 55.78 41.44 44.06 Self-Refine (Madaan et al., 2023) 41.58 49.56 42.75 29.64 49.65 57.62 54.73 43.24 45.19 S3 Agent (Wang et al., 2024) 42.51 51.52 41.34 31.13 48.85 61.44 56.31 38.96 45.34 DeepSeek-V3 (Liu et al., 2024) Direct (Liu et al., 2024) 56.27 65.50 58.65 41.61 62.06 73.72 69.47 52.70 58.86 Zero-shot CoT (Kojima et al., 2022) 56.18 63.97 59.71 42.51 61.83 71.18 70.52 53.15 58.71 Plan-and-Solve (Wang et al., 2023) 55.16 63.97 56.18 42.21 63.33 70.76 71.05 56.08 58.81 Self-Refine (Madaan et al., 2023) 53.86 61.79 55.83 41.31 59.54 69.06 64.73 52.25 56.40 S3 Agent (Wang et al., 2024) 54.69 61.57 56.89 43.41 59.42 71.18 66.31 50.22 56.81 DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) Direct (Guo et al., 2025) 50.41 60.48 50.17 36.82 57.12 65.25 64.73 49.54 53.41 Zero-shot CoT (Kojima et al., 2022) 51.25 60.91 51.94 37.42 56.55 68.64 63.15 47.29 53.62 Plan-and-Solve (Wang et al., 2023) 51.16 60.91 50.53 38.32 55.86 66.52 62.63 47.29 53.26 Self-Refine (Madaan et al., 2023) 44.65 52.18 47.70 31.73 52.06 57.20 60.00 45.94 47.96 S3 Agent (Wang et al., 2024) 51.25 59.60 51.94 39.22 58.04 69.49 65.26 48.87 54.29 Qwen-3-14B (Yang et al., 2025) Direct (Yang et al., 2025) 54.88 66.59 58.65 35.62 59.08 72.03 70.52 47.29 56.76 Zero-shot CoT (Kojima et al., 2022) 55.72 68.34 56.18 38.02 59.19 69.91 70.52 48.64 57.27 Plan-and-Solve (Wang et al., 2023) 54.97 66.81 58.65 35.62 57.81 71.18 68.94 45.94 56.24 Self-Refine (Madaan et al., 2023) 54.41 65.93 56.89 36.52 58.27 71.61 66.84 47.52 56.14 S3 Agent (Wang et al., 2024) 54.13 65.06 57.24 36.52 58.50 71.18 67.36 47.97 56.09 O4-mini (OpenAI, 2025) Direct (OpenAI, 2025) 61.30 70.08 65.01 46.10 70.22 77.54 75.78 63.96 65.29 Zero-shot CoT (Kojima et al., 2022) 62.60 69.86 64.66 50.89 71.37 77.11 77.89 65.54 66.52 Plan-and-Solve (Wang et al., 2023) 61.67 69.86 64.31 48.20 72.41 77.96 77.89 67.11 66.47 Self-Refine (Madaan et al., 2023) 61.20 69.21 60.77 50.59 70.34 75.00 74.73 65.99 65.29 S3 Agent (Wang et al., 2024) 61.76 69.65 62.19 50.59 71.49 76.69 75.78 66.89 66.11
@@ 4.1 Experimental Settings
We use the standard financial question answering benchmark FAMMA (Xue et al., 2024) for experiments. FAMMA is a benchmark for financial multilingual multi-modal question answering, which includes English, Chinese, and French. Since current RLLMs are unimodal and cannot process multimodal information, we utilize the Basic Txt dataset in FAMMA, which converts multimodal data into textual information using OCR. The final dataset used for evaluation comprises 1,945 entries.
We conduct experiments on seven backbones, including five LLMs and three RLLMs: Meta-Llama-3.1-8B-Instruct (Grattafiori et al., 2024), GPT-4o-mini (Hurst et al., 2024), Gemini-1.5-flash (Team et al., 2024), Qwen-2.5-32B (Yang et al., 2024), DeepSeek-V3 (Liu et al., 2024), DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025), thing mode of Qwen-3-14B (Yang et al., 2025), O4-mini (OpenAI, 2025). Detailed settings can be found in the Appendix LABEL:appendix:details.
@@ 4.2 Analysis of Prompting Methods
Larger LLMs exhibit increased robustness to various prompting methods. It is evident that for larger LLMs, such as Gemini-1.5-flash, Qwen-2.5-32B, and DeepSeek-V3, the performance gap between different prompting methods is not significant, with performance fluctuations remaining under 2%. This indicates that LLMs with superior performance are less affected by input prompts, demonstrating greater robustness to variations in input.
Performance Direct Translate-en (Shi et al., 2023) Self-Refine (Madaan et al., 2023) S3 Agent (Wang et al., 2024) CLP (Qin et al., 2023) Meta-Llama-3.1-8B-Instruct (Grattafiori et al., 2024) Chinese 16.54 16.54 (+0.00) 18.97 (+2.43) 19.22 (+2.68) 23.35 (+6.81) French 15.38 16.66 (+1.28) 13.46 (-1.92) 21.15 (+5.77) 17.30 (+1.92) Overall (Chinese+French) 16.22 16.57 (+0.35) 17.46 (+1.24) 19.75 (+3.53) 21.69 (+5.47) GPT-4o-mini (Hurst et al., 2024) Chinese 33.09 34.54 (+1.45) 34.79 (+1.70) 31.63 (-1.46) 38.19 (+5.10) French 28.84 30.76 (+1.92) 28.84 (+0.00) 32.69 (+3.85) 32.05 (+3.21) Overall (Chinese+French) 31.92 33.50 (+1.58) 33.15 (+1.23) 31.92 (+0.00) 36.50 (+4.58) Gemini-1.5-flash (Team et al., 2024) Chinese 40.38 38.68 (-1.70) 36.73 (-3.65) 39.65 (-0.73) 37.22 (-3.16) French 27.56 29.48 (+1.92) 26.92 (-0.64) 27.56 (+0.00) 29.48 (+1.92) Overall (Chinese+French) 36.86 36.15 (-0.71) 34.03 (-2.83) 36.33 (-0.53) 35.09 (-1.77) Qwen-2.5-32B (Yang et al., 2024) Chinese 41.84 44.52 (+2.68) 44.28 (+2.44) 42.82 (+0.98) 43.30 (+1.46) French 29.48 29.48 (+0.00) 33.97 (+4.49) 35.89 (+6.41) 35.25 (+5.77) Overall (Chinese+French) 38.44 40.38 (+1.94) 41.44 (+3.00) 40.91 (+2.47) 41.09 (+2.65) DeepSeek-V3 (Liu et al., 2024) Chinese 57.42 58.63 (+1.21) 55.96 (-1.46) 57.66 (+0.24) 61.31 (+3.89) French 50.00 50.00 (+0.00) 47.43 (-2.57) 44.23 (-5.77) 57.69 (+7.69) Overall (Chinese+French) 55.37 56.26 (+0.89) 53.61 (-1.76) 53.96 (-1.41) 60.31 (+4.94) DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) Chinese 48.41 49.87 (+1.46) 46.47 (-1.94) 52.06 (+3.25) 50.85 (+2.44) French 49.35 42.94 (-6.41) 44.87 (-1.16) 49.35 (+0.00) 52.56 (+3.21) Overall (Chinese+French) 48.67 47.97 (-0.70) 46.03 (-2.64) 51.32 (+2.65) 51.32 (+2.65) Qwen-3-14B (Yang et al., 2025) Chinese 57.66 56.69 (-0.97) 55.23 (-2.43) 55.47 (-2.19) 53.77 (-3.89) French 48.71 50.00 (+1.29) 42.3 (-6.41) 48.71 (+0.00) 51.28 (+2.57) Overall (Chinese+French) 55.20 54.85 (-0.35) 51.67 (-3.53) 53.61 (-1.59) 53.08 (-2.12) O4-mini (OpenAI, 2025) Chinese 65.45 65.93 (+0.48) 64.23 (-1.22) 67.63 (+2.18) 64.54 (-0.91) French 62.17 61.53 (-0.64) 60.25 (-1.92) 61.53 (-0.64) 63.46 (+1.29) Overall (Chinese+French) 64.55 64.72 (+0.17) 63.13 (-1.42) 65.69 (+1.14) 64.24 (-0.31)